OpenStudy (anonymous):

f(x)=3x-7 and g(x)=-2x-6. Find (f times g)(4)

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=3x-7 }\) \(\large\color{red}{ g(x)=-2x-6 }\) \(\large\color{slate}{ (\color{blue}{ f }\times \color{red}{ g })(x)=\color{blue}{ (3x-7) }\times \color{red}{ (-2x-6)}}\)

OpenStudy (solomonzelman):

you can go ahead without expanding anything, \(\large\color{slate}{ (\color{blue}{ f }\times \color{red}{ g })(x)=\color{blue}{ (3x-7) }\times \color{red}{ (-2x-6)}}\) \(\large\color{slate}{ (\color{blue}{ f }\times \color{red}{ g })(\color{green}{4})=\color{blue}{ (3\color{green}{(4)}-7) }\times \color{red}{ (-2\color{green}{(4)}-6)}}\)

OpenStudy (solomonzelman):

simplify the last expression and that would be the answer

OpenStudy (anonymous):

0?

OpenStudy (solomonzelman):

how are you getting zero? is any of the terms on the product equal to zero?

OpenStudy (solomonzelman):

terms **in the product

OpenStudy (anonymous):

240?